A construction company has adjoined a 5000 ft^2 rectangular enclosure to its office building. Three sides of the enclosure are fenced in. The side of the building adjacent to the enclosure is 500 ft long and a portion of this side is used as the fourth side of the enclosure. Let x and y be the dimensions of the enclosure, where x is measured parallel to the building, and let L be the length of fencing required for those dimensions.

(a) Find a formula for L in terms of x and y.
L(x,y)=
b) Find a formula that expresses L as a function of x alone.
L(x)=
(c) What is the domain of the function in part (b)? Express as an interval.

Domain =
1
Expert's answer
2020-11-27T14:58:49-0500

"\text{The length of the fence is the perimeter of the rectangle minus the common side}"

"\text{with the office building}"

"a) L(x,y) =2*(x+y)-x = 2y+x"

"b) S =x*y; y = \frac{S}{x}= \frac{5000}{x}"

"L(x) = \frac{1000}{x}+x"

"c) \text{Since the length is positive and the side parallel to the building and the office}"

"\text{ cannot exceed 500 ft}"

"\text{Domain} =(0,500]"

Answer: "L(x,y) = 2y+x" ; "L(x) = \frac{1000}{x}+x" ;"\text{Domain} =(0,500]"